Week of 200206

On Call: Sophie/James

Monday:

Log: Nothing

Tuesday:

Log: Nothing

Wednesday

Log: Nothing

Thursday

Log: Expired hostcerts on half of the oplapro machines. Removed from castorgridsc cluster until they can be re-issued.

Actions:

Discussion: Planning new sensor to check for this.

Friday

Log:

Actions:

Discussion:

-- GavinMcCance - 23 Feb 2006

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-02-23 - GavinMcCance
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback